# Towards Clean-Label Backdoor Attacks in the Physical World

### v2

@@ Appendix A Natural Physical Backdoors
Appendix B Dataset
B.1 Ethics & Data Protection
@@ B.4 Dataset Structure
False triggers. To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix C.4), we select three additional false triggers for each of the main triggers. Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker). We also select objects that have similar locations on the face as false triggers (e.g., headbands are selected as false triggers for the hat trigger and the sunglasses trigger). In Figure 17, each row corresponds to a main trigger on the left, followed by three false triggers on the right.
@@
B.5 Animal Classification Dataset
@@
Appendix C Experiment Settings
C.1 Computational Resources.
All the experiments are conducted on the two servers. The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one contains 6 RTX A5000 24 GB GPUS. The code is implemented in Pytorch. We develop the framework based on the previous works [49, 15]. We use Pytorch as the framework for our model training.
C.2 Training Settings.
@@ Appendix C Experiment Settings
Appendix D Additional Experiments
D.1 Blended Label-Consistent Attack
@@
D.2 Experiments on Different Models
@@
D.3 Scaling the Dataset Size
@@
D.4 Evaluating Digital Trigger for CLPBA
@@
D.5 Evaluating transferability of CLPBA
@@
Appendix E Additional Visualization
E.1 Poisoned samples of different poisoning algorithms
@@
As shown in Figure 12, the perturbations of GM appear to reconstruct the sunglasses trigger. In FM, the perturbations are more uniform and fine-grained, but in some poisoned samples the perturbations resemble those crafted by GM (the image in the last column). We believe that the reconstruction of the sunglasses trigger is the key to backdoor the victim model. However, not all samples exhibit the perturbation pattern of a trigger. GM is more successful than FM because it successfully reconstructs the trigger in the perturbation in most of the samples.
E.2 Additional imperceptibility comparison of regularizations
@@
Appendix F Defenses
@@
F.1 Evaluation on filtering defenses.
@@
F.2 Evaluation on firewall defenses.
@@
Appendix G Limitations and future works
@@
Appendix H Societal impact

### v3

@@ 3 Natural Physical Backdoors
4 Clean-Label Physical Backdoor Attack
4.1 Problem Formulation
The effectiveness of the Naive attack, as introduced in Section 3, is highly dependent on the combination of source class, target class, and physical trigger and thus may not generalize to realistic settings. Furthermore, the existence of a physical trigger in inserted poisoned images may raise suspicion. In this section, we introduce the problem formulation and poisoning algorithms of CLPBA, which crafts triggerless poisoned samples that indirectly plant the backdoor (activated by a physical trigger) into the models.
where $C$ is the constraint set of the perturbations: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$. $\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class. The poison set $P$ is a small subset of $D_{t}$. Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples. The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples. We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data. Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set. With the above formulation, we introduce three poisoning algorithms for CLPBA based on existing clean-label data-poisoning methods:
  • Feature Matching (FM) [41]: This method crafts perturbations that shift poisoned samples closer to source trigger distribution $\widetilde{\mathcal{D}}_{s}$ in the feature space, thus representing them as dirty-label samples with trigger:
    $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (5)
  • Gradient Matching (GM) [46]: This method crafts perturbations that align the gradient of the model on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
4.2 Threat Model
Attacker’s capability. In our threat model, the training dataset is collected from multiple sources, and there exists an attacker (e.g., data contributors, data collectors) who can access and perturb a small portion of the data to manipulate the behavior of the victim model. Following previous works [41, 46], we assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process. The attack is one-to-one: any instances from the source class are misclassified as belonging to the target class with the presence of the trigger. For example, the source class is an employee in a company who can wear a special pair of sunglasses to fool a facial recognition system into classifying him as the CEO so that he could get an illegal privilege, such as access to confidential documents or high-ranking facilities. Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
Figure 5: Physical triggers in the collected datasets
4.3 Data Collection
Table 3: ASR (%) of poisoning algorithms on selected source-target pairs, averaged over 10 runs with standard deviation in brackets.
Scenario Poison ratio $\alpha$ Naive MP FM GM
(a) Facial Recognition (sunglasses) trigger
finetuning 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
transfer 5% 0.0 (0.0) 2.2 (0.5) 1.8 (1.3) 81.8 (5.6) 77.8 (9.6) 91.2 (7.6) 92.5 (8.5)
10% 0.0 (0.0) 2.4 (1.8) 3.6 (1.2) 79.3 (2.3) 64.5 (25.7) 93.6 (6.9) 97.2 (2.8)
(b) Animal Classification (tennis) trigger
fine-tuning 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
transfer 5% N/A 1.1 (0.7) 3.2 (2.5) 34.5 (1.9) 29.7 (3.0) 62.0 (2.1) 75.6 (5.5)
10% N/A 1.5 (1.2) 2.8 (1.2) 33.6 (2.8) 34.3 (2.3) 92.8 (1.2) 88.4 (3.7)
5 Experiments
5.1 Experiment Settings
We split our datasets with a ratio of 8:2 for train and test data. We use 50% of source-class trigger images to craft poisoned samples, and the remaining 50% is used to evaluate ASR. Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper. We choose ResNet50 [18], pre-trained on the VGGFace2 dataset [4], to be the victim model for Facial Classification, while ResNet18 pre-trained on Imagenet [40] is used For Animal classification. We evaluate the attacks under transfer learning and fine-tuning settings, two common settings in clean-label data poisoning [41, 46]. For transfer learning, the feature extractor is frozen and only the last linear layer is trained, while for fine-tuning, we set a low learning rate of 0.001 to train the victim model on the poisoned dataset. To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters. For the Facial Classification task, we select the poison-pair $9-5$ (as class labels in our dataset) for our experiment. This poison pair is manually selected to avoid dataset bias, as explained in Section 3. For the Animal Classification task, we select the poison-pair cat-fox.
5.2 Baseline evaluation
From the experiment results demonstrated in Table 3, we have the following observations:
  • GM is the most effective poisoning algorithm. GM is especially effective for the fine-tuning scenario, which is a harder scenario compared to transfer learning since the weights of all layers of the victim model are updated. Matching the gradient update of $L_{adv}$ allows the GM perturbations to utilize the backward information of all layers, instead of only the last linear layer as in FM. MP is not effective because the poisoned samples overfit the updated parameters of the unrolled surrogate model in the inner loop, which greatly differ from the victim model.
  • Impact of $\epsilon$ and $\alpha$. While increasing either $\epsilon$ or $\alpha$ would improve the performance of attacks, we observe that $\epsilon$ plays a more important role than $\alpha$, especially for fine-tuning scenario. A high value of $\epsilon$ is critical for the poisoning objective to converge, while a high value of $\alpha$ does not guarantee that every poisoned sample is effective.
  • Effectiveness of CLPBA varies across datasets. As shown in Table 3, CLPBA is significantly less effective in Animal Classification than in Facial Recognition. This reduced performance is due to the tennis trigger being as dominant as the sunglasses trigger while also appearing in varying locations across the trigger samples, making it harder for the model to learn.
5.3 Analysis of CLPBA
Existence of adversarial features. To justify the existence of adversarial features, we take inspiration from [24] to test whether the model trained on poisoned target samples could successfully predict source-class samples. We design the experiment as illustrated in Figure 6. Then, we filter out all source-class samples before relabelling all GM-poisoned samples to the source class. We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously. To avoid randomness, we set a prediction threshold of 85% after the softmax function. Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples. This shows that the perturbations on poisoned target instances contain adversarial features representative of the source-class distribution. To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [32] as the source class, and the accuracy is 0%. The setup of this experiment can be found in Appendix D.
(a) Clean model
(b) Poisoned model
Figure 7: Feature space visualization.
6 Improving the Effectiveness of CLPBA
6.1 Pixel regularization
where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$. $M$ encourages the perturbation to align with the underlying image content. In Equation 12, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$. The weighting factor $\alpha$ is set to 5. Both terms in Equation 11 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels.
Side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 9 and Figure 10. Our pixel regularization leads to a high ASR even for small values of $\epsilon=[4,8]$, while the ASR for standard $\ell_{\infty}$ regularization is below 20%. At the same time, our pixel regularization also has higher PSNR values for the poisoned images at every $\epsilon$, indicating better perceptual stealthiness. Table 4 demonstrates the improved effectiveness of pixel regularization across different poison pairs.
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 9: Visualization between standard $\ell_{\infty}$-regularization and our Pixel Regularization. Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=16/255$)
Table 4: ASR (%) of GM attack + tennis with two regularizers.
Regularizer Fox Kangaroo Mongoose Sea-lion Tapir
$\ell_{\infty}$ regu 9.6 (4.1) 8.4 (0.8) 5.3 (0.4) 13.4 (2.9) 7.8 (5.0)
pix. regu 30.9 (10.4) 12.5 (1.6) 4.4 (1.0) 24.2 (2.3) 26.6 (5.2)
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 10: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization on different values of $\epsilon$. Higher PSNR indicates better perceptual quality of poisoned images.
6.2 Ensembling and Retraining
Following [46], we also apply ensembling and retraining for CLPBA methods in the fine-tuning setting of the Animal Classification task. In our experiment, we apply an ensemble of 3 ResNet18 models; each of these models will retrain on the poisoned dataset after every 200 attack iterations. The attacker optimizes poisoned samples with GM for 1000 iterations. Table 5 shows that applying these techniques significantly improves the attack success rate.
Table 5: ASR (%) of GM ($\alpha=0.1,\epsilon=16/255$) on the Animal Classification task with tennis trigger.
Method ASR (%)
Baseline 17.8 (7.7)
Retraining 25.5 (4.2)
Ensemble+Retraining 29.7 (3.1)
7 Defending against CLPBA
Filtering Defenses
Metrics AC[3] SS[49] SPECTRE[17] SCAN[48]
ER (%) 62.16 97.30 72.97 0
SR (%) 15.77 19.77 19.50 0
ASR (%) 99.7 0.0 99.7 99.7
Firewall Defenses
Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[56]
TPR (%) 0.0 79.6 0.0 0.0
FPR (%) 6.7 65.3 16.7 0.9
8 Accidental Backdoor Activations
Poison pair Trigger Loss type ASR ACC FPR SR
9-5 Fake beard Original 98.7 99.9 50.8 12.5
With repel 98.7 99.9 49.7 3.9
7-3 Sunglasses Original 97.5 99.3 0.7 13.9
With repel 97.5 99.4 0.9 2.5
[…]
One limitation of using the repel terms is that the attacker must possess trigger samples from non-source classes. To circumvent this, the attacker can patch clean images with digital triggers, mapping them to appropriate facial areas.
9 Conclusion
Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks. We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances. Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples. A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness. To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class.
Supplementary Material
Appendix A Dataset
A.1 Ethics & Data Protection
@@ A.4 Dataset Structure
False triggers. To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix B.4), we select three additional false triggers for each of the main triggers. Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker). We also select objects that have similar locations on the face as false triggers (e.g., headbands are selected as false triggers for the hat trigger and the sunglasses trigger). In Figure 20, each row corresponds to a main trigger on the left, followed by three false triggers on the right.
@@
A.5 Animal Classification Dataset
@@
Appendix B Experiment Settings
B.1 Computational Resources.
All the experiments are conducted on the two servers. The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one has 6 RTX A5000 24 GB GPUS. The code is implemented in Pytorch. We develop the framework based on the previous works [46, 15]. We use Pytorch as the framework for our model training.
B.2 Training Settings.
@@ Appendix B Experiment Settings
Appendix C Additional Experiments
C.1 Blended Label-Consistent Attack
@@
C.2 Experiments on Different Models
@@
C.3 Scaling the Dataset Size
@@
C.4 Evaluating Digital Trigger for CLPBA
@@
C.5 Evaluating transferability of CLPBA
@@
Appendix D Additional Visualization
@@
Visualization of poisoned samples. As shown in Figure 13, the perturbations of GM appear to reconstruct the sunglasses trigger. In FM, the perturbations are more uniform and fine-grained, but in some poisoned samples the perturbations resemble those crafted by GM (the image in the last column). We believe that the reconstruction of the sunglasses trigger is the key to backdoor the victim model. However, not all samples exhibit the perturbation pattern of a trigger. GM is more successful than FM because it successfully reconstructs the trigger in the perturbation in most of the samples.
@@
Figure 15: Visualization of poisoned squirrel images with the standard $\ell_{\infty}$ regularization
Figure 16: Visualization of poisoned squirrel images with pixel regularization
Appendix E Defenses
@@
E.1 Evaluation on filtering defenses.
@@
E.2 Evaluation on firewall defenses.
@@
Appendix F Limitations and future works
@@
Appendix G Societal impact